# Two Sides of the Same Coin: Large-scale Measurements of Builder and Rollup after EIP-4844

### v2

@@ Abstract.
Web3 is reshaping decentralized ecosystems through innovations like Ethereum. Recently, EIP-4844 is implemented in Ethereum to support its Layer-2 scaling solutions, which introduces a new 128 KB data structure called blob. This upgrade incorporates type-3 transactions with blobs to verify data availability and reduce gas costs for rollups, significantly affecting the strategies of both builders and rollups. In this paper, we present an in-depth study of emerging strategies in builder and rollup markets after EIP-4844, containing hundred million transactions. We find that the efficiency of builder and rollup strategies is interdependent, akin to two sides of the same coin—both cannot be optimized simultaneously. That is, when builders operate efficiently, rollups tend to overpay in fees, conversely, when rollups optimize their costs, builders may incur losses in inefficient transaction selection. From the side of builders, our results show that 29.48% of these blocks have been constructed inefficiently, which does not produce sufficient profits for builders. Through our evaluation from the side of rollups, we find that over 72.53% of type-3 transactions pay unnecessary fees, leading to notable economic costs of rollups. Our work provides critical insights into optimizing block construction and transaction strategies, advancing the economic efficiency and data scalability of Web3 infrastructures, yet, much like balancing a seesaw, the efficiency of builders and rollups cannot be optimized concurrently.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the Conference 2025; June 28–October 02, 2018; NY, USAisbn: 978-x-4503-XXXX-X/18/06ccs: General and reference Empirical studiesccs: General and reference Measurementccs: General and reference Evaluation
@@ 1. Introduction
Web3 leverages decentralized technologies such as blockchain to enable peer-to-peer interactions without intermediaries. A groundbreaking and transformative application of decentralized technology is Ethereum, an open and decentralized blockchain platform that facilitates the development and execution of smart contracts and decentralized applications (Gilbert, 2022; Wang et al., 2022b).
In the Ethereum mainnet, the Proposer Builder Separation (PBS) mechanism was proposed to separate the role of the miner by introducing the builder and the proposer (Buterin, 2021a). The task of builders is to construct the block, where they sort the transactions to maximize the extractable value and submit the block to proposers (Daian et al., 2020; Heimbach and Wattenhofer, 2022). Each transaction within the block originates from two sources: the public mempool, which is accessible to everyone, and private channels that deliver transactions directly to specific builders (Qin et al., 2022).
Ethereum Layer-2 refers to scaling solutions built on top of the Ethereum mainnet to enhance its performance and throughput, with rollups being one of the most prominent approaches that bundle multiple transactions into a single batch before submitting them to the mainnet. Ethereum has implemented EIP-4844 in the Dencun upgrade to improve network scalability and lower transaction costs (Ethereum.org, 2024a). This upgrade presents blob, a new data format of 128 KB designed to facilitate rollups on Ethereum’s Layer-2, alongside blob-carrying transactions known as type-3 transactions (Vitalik Buterin, 2023). By offloading data storage to blobs, the gas fee is significantly reduced for rollups. Type-3 transactions are much larger and therefore take more space in blocks (Park et al., 2024; Gomez, 2024b). Blocks with large size take longer to propagate through the network. On the one side, builders need to balance the block size with the profit they can obtain (Wahrstätter, 2024b; Gomez, 2024a). They also need to decide whether to choose type-3 transactions in their blocks. This results in different behaviors among builders in response to type-3 transactions. On the other side, rollups need to issue their blobs into type-3 transactions and wait for being included in blocks. Type-3 transactions always carry the different number of blobs and different fees, which indicate various strategies of rollups (Cui, 2024). Evidently, the strategies among rollups are also complicated.
We evaluate the strategies of builders and rollups after EIP-4844. We find that the efficiency of their strategies is like the two sides of the same coin. When the builders strategy is efficient, rollups may need to pay more fees. Oppositely, when rollups adopt the efficient strategy and offer lower fees, builders may face block profits lower than potential profits. In the side of builders, we first examine the new strategy that builders employed after EIP-4844 (Section 4). Although the average total block size has been increasing to nearly 400 KB, the size of transactions from public mempool in the block has decreased from 150 KB before the upgrade to about 30 KB. Builders have taken to shrinking the size of the rest of the block to contain type-3 transactions, which prompts us to delve deeper into the efficiency of this strategy. We recognize that the builder strategy is not efficient enough after EIP-4844. In the side of rollups, they need to pay higher fees to ensure inclusion in efficient blocks. However, we find that most type-3 transactions issued by rollups pay too high fees (Section 5). Our primary contributions are:
@@
    We build a large-scale dataset to quantify builders and rollups market in Ethereum from after EIP-4844 to August, which includes 319,529,950 transactions with 1,336,822 type-3 transactions. To the best of our knowledge, we are the first to illuminate different strategies and efficiencies of builders and rollups in the network after EIP-4844. Our dataset provides the most complete labelling of builders and rollups so far.
@@
    We identify the strategies that different builders adopt after EIP-4844. By examining the size of transactions versus the profit paid to builders, we provide a model for measuring efficient blocks for builders. We find that 29.48% of the blocks containing type-3 transactions do not give the builder enough profits to compensate for the loss of excluding other transactions. This will assist builders in improving their strategies for constructing blocks for higher profit margin.
@@
    We explore the sending patterns of different rollups for blob and type-3 transactions and find that the behavior within the market is distinctive. We measure efficient type-3 transactions for rollups and observe that 72.53% of type-3 transactions in the market are inefficient. This work contributes to how rollups can optimize their pricing of type-3 transactions to achieve cost savings.
  4. (4)
    Furthermore, we evaluate the losses of other flawed strategies in the market and reveal significant losses of 186.92 ETH for rollups. We also characterize the impact of one misuse of blobs. It induces a delay of over 19 seconds in the type-3 transactions of rollups.
2. Backgrounds
@@
In this section, we provide an overview of Ethereum builders, rollups and EIP-4844.
2.1. Ethereum
Ethereum stands as a leading decentralized blockchain platform (Buterin et al., 2013). With an upgrade known as the Merge (Foundation., 2023), Ethereum has transitioned from a PoW consensus, which is energy intensive, to a PoS mechanism. The Proposer Builder Separation (PBS) mechanism has been proposed as an important development after the Merge (Ethereum.org, 2023). The traditional role of validators, formerly known as miners, is divided into separate entities: block builders and block proposers (i.e., the validators themselves) (Buterin, 2021a, b). Flashbots introduced MEV-Boost as an intermediate measure to enable the implementation of PBS (Wahrstätter, 2023). The other roles are searchers and relays. The role of searchers is to identify and leverage profit extraction opportunities in the public mempool (Qin et al., 2022; Wang et al., 2022a; Zhou et al., 2023). Once they finish extracting the value, they would send the transaction bundles to builders via private channels (Mamageishvili et al., 2024). Relays are introduced to establish the trust required between block builders and proposers (Flashbots, 2022).
As shown in Figure 1, builders receive two types of transactions, public transactions from public mempool (Wackerow, 2024a) and private transactions from private channels (Raun et al., 2023). In addition to the transaction priority fee and coinbase transafer, the builder may profit by collaborating privately with the transaction sender (EigenPhi, 2024). Builders prioritize transactions with higher priority fee and subsequently rank transactions by their potential revenue, before determining which transactions to include in blocks. This creates a block space race in which users can bid for faster inclusion through higher priority fees (Liu et al., 2022). Users who want faster confirmations can offer higher priority fees, while those willing to wait can pay just the base fee.
Currently, it is optional for Ethereum validators to participate in PBS (Wahrstätter, 2024b). Validators who opt not to use PBS and instead build blocks independently are known as Vanilla builders. Vanilla builders receive no private order flow because they are not specialized in block building tasks. They complete both block building and validation tasks simultaneously and currently account for 10% of all builders in the network (Wahrstätter, 2023).
2.2. Rollup
Rollups are Layer-2 scaling solutions designed to improve the efficiency and throughput of the Ethereum network by offloading transaction processing from the Ethereum mainnet. To protect the data availability of Ethereum (Pomerantz, 2024), rollups work by bundling or rolling multiple transactions into a single batch. After processing, the rollup smart contract on Ethereum updates the state with a single, compressed transaction that includes the aggregated data of all individual transactions. It will be submitted to the mainnet, thereby reducing the computational load and gas fee for each individual transaction. Rollups maintain the security and decentralization of Ethereum while providing faster and cheaper transactions, making them a crucial component of the Ethereum scalability roadmap (Smith, 2024).
@@ 2.2. Rollup
2.3. EIP-4844
@@
Rollups used calldata to store state roots and proofs (Pomerantz, 2024), ensuring that the rollup state could be re-established even if a transaction was challenged. Calldata is the payload contained in a transaction that can be used to store arbitrary information on the chain. While calldata is cost-effective for smaller amounts of data, it is permanently stored in Ethereum state and is therefore relatively expensive for larger data requirements. The high cost of calldata storage becomes a bottleneck for scaling.
@@
In order to improve the scalability and data availability of the network, Ethereum proposed a new sharding design, Danksharding (Ethereum.org, 2024b). Danksharding is a unified sharding design that differs from traditional sharding by managing data storage and availability in a more streamlined way. Instead of increasing the space for transactions, Ethereum sharding provides more space for blobs of data, which the Ethereum protocol itself does not interpret. EIP-4844, also known as Proto-Danksharding, implements the same transaction formats and validation rules as the new framework and logic of Danksharding. EIP-4844 was launched in March as a core part of the Dencun upgrade in 2024 (Ethereum.org, 2024a), thereby facilitating the network’s preparation for full Danksharding by allowing the inclusion of large data in transactions (Pomerantz, 2024). The EIP-4844 represents a significant milestone in Ethereum’s roadmap (Ethereum.org, 2024c), aiming to improve the scalability, security, and usability of the network.
@@
There are three types of transaction before EIP-4844, type-0, type-1 and type-2 transaction, and type-2 transactions always bring priority fee for inclusion by builders quickly. EIP-4844 focuses on type-3 transactions (Park et al., 2024) which are similar to regular transactions like type-2 transactions, but carry an additional data component called ‘blob’. These oversized blobs (128 KB) are much cheaper than the equivalent amount of calldata because they are not permanently stored on the Ethereum (Pomerantz, 2024), and rollups can also send their type-3 transactions from public mempool or private channels. This new type of transaction meets the growing need for efficient data processing. The blob data remains inaccessible during the execution of the Ethereum virtual machine, and the validator can only see the commitment to the blob, but these blobs can be used for Layer-2 scaling solutions and other off-chain protocols that require large amounts of data throughput (Gogol et al., 2024a).
3. Dataset

### v3

# The Price of Cheaper Data: Measuring the Strategic Inefficiencies in the Post-EIP-4844 Ethereum Market

@@ Abstract.
High transaction fees on Ethereum have long hindered its scalability and user adoption. The recent Ethereum EIP-4844 upgrade aims to alleviate the scalability issue by introducing the “blob”, a new data structure for Layer-2 rollups. Instead of using expensive blockchain storage, blobs provide a cheaper, separate data layer with its own fee market, which drastically lowers data availability costs. This change, while lowering transaction costs, has created a new high-stakes economic game for block builders and rollups. However, the dynamics of this game remain poorly understood. In this paper, we conduct the first large-scale empirical analysis of the post-EIP-4844 ecosystem, leveraging a dataset of $319.5$ million transactions, including $1.3$ million blob-carrying type-3 transactions. Our analysis demonstrates clear evidence of a structural shift towards the utilization of cheap blobs over expensive transactions for Layer-2 data posting: while average block size grew $2.5\times$, the space consumed by expensive transactions in the public mempool shrank from over 150 KB to just 30 KB. Yet, this scalability success masks widespread economic inefficiencies on both sides of the market. On the builder side, $29.48\%$ of blocks containing blobs are constructed sub-optimally, earning less revenue than that could be achieved otherwise. On the rollup side, we identify that flawed submission strategies have led to $186.92$ ETH in direct losses and average inclusion delays of over $19$ seconds. Moving beyond characterization, our work offers actionable solutions. For builders, we develop an optimal pricing model derived from a formal first-price auction framework, allowing builders to make provably profitable inclusion decisions. For rollups, we prove that batching multiple blobs into a single transaction is a dominant, utility-maximizing strategy. Our work presents the first data-driven characterization of the post-EIP4844 market, revealing critical insights for the future development of Ethereum’s scalability.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: General and reference Empirical studiesccs: General and reference Measurementccs: General and reference Evaluation
@@ 1. Introduction
1.1. Background
Ethereum is a decentralized, blockchain-based smart contract platform. The ability to execute code trustlessly allows Ethereum to serve as a foundational layer for a wide array of decentralized applications, including decentralized Finance (DeFi), decentralized voting systems, and stablecoins. Executing transactions and smart contracts on the network consumes computational resources, which requires a fee to manage demand and incentivize network participants. This fee is calculated using“Gas”—a unit that measures the complexity of an operation—and is paid in the platform’s native cryptocurrency, Ether (ETH). This economic model has proven highly effective, fueling the platform’s widespread adoption. Data from Glassnode indicates that Ethereum has become one of the world’s leading blockchains, with its total address count reaching $250$ million and surpassing Bitcoin in several key on-chain activity metrics (Glassnode, 2025).
As the Ethereum user base has grown, the mainnet has faced significant network congestion, leading to limited throughput, extended transaction confirmation times, and prohibitively high fees. To address these critical scalability limitations, the community has championed external solutions known as Layer-2 protocols. Among the Layer-2 protocols, Rollups have emerged as a prominent approach. Rollup solutions enhance efficiency by bundling hundreds of transactions into a single batch on a separate chain and then submitting a compressed summary of this batch to the Ethereum mainnet for final settlement. By offloading the bulk of computation and execution in this manner, Rollups drastically reduce congestion and lower costs for users, all while inheriting the robust security guarantees of the mainnet.
The growing adoption of Layer-2 solutions, particularly Rollups, placed significant economic pressure on their operators. To ensure data availability and security, these protocols were required to post compressed transaction data to the expensive “Calldata” space on the Ethereum mainnet—used for verifying data in Layer-2 transactions As user activity scaled, the frequency and cost of these “Calldata” submissions increased commensurately, creating a major operational burden. At its peak, this “data tax” on Calldata often cost major Rollups a collective sum exceeding $10,000$ USD per day, according to analyses on Dune Analytics (0xRob, 2025). To alleviate this, the Ethereum Dencun upgrade introduced EIP-4844, which created a dedicated and more cost-effective data space for Rollups known as “blobs”. Accessed via a new Type-3 transaction, blobs allow Rollups to offload data from Calldata, thereby significantly reducing their operational costs (Ethereum.org, 2024a; Buterin et al., 2023). This new dynamic structurally alters the decision-making for both block builders and rollup operators. Block builders, whose primary task is to construct the most profitable block by sorting transactions (Daian et al., 2020; Heimbach and Wattenhofer, 2022), now face a critical dilemma. Including large, blob-carrying transactions may increase fee revenue, but it also increases the block’s propagation time, heightening the risk of it being orphaned and losing all profit. Concurrently, rollup operators must compete to have their Type-3 transactions included by these discerning builders. This has led to multifaceted strategies (Crapis et al., 2024), where Rollups must decide on optimal blob bundling and offer compelling priority fees to incentivize builders to accept the added risk. The result is a new, complex marketplace for blob inclusion with diverse bidding behaviors (Cui, 2024).
Failing to properly understand this market can have adverse consequences. For builders, a persistent reliance on suboptimal block construction strategies may prevent them from capturing the revenue opportunities provided by type-3 transactions. Over time, this revenue gap erodes their competitiveness relative to faster-adapting peers, ultimately forcing them out of the builder market (Wang et al., 2025). For rollups, inefficient blob-carrying strategies result in consistently higher inclusion costs and longer delays, a burden that is inevitably passed on to the users of their Layer-2 networks.
1.2. Contribution
Our research is driven by the following three research questions, where efficiency is understood in terms of profitability.
  • RQ1: What are current builder strategies for block construction, and are they efficient?
  • RQ2: What are current rollup strategies for submitting blobs, and are they efficient?
  • RQ3: What is the profit-optimal strategy for builders and rollups?
To answer the questions and illuminate the unexplored dynamics of the post-EIP-4844 market, we begin by constructing the first large-scale dataset dedicated to analyzing the strategic interactions between builders and rollups in this new era. This comprehensive dataset, encompassing over $319$ million transactions—including $1.3$ million Type-3 transactions, serves as the bedrock for our entire investigation. We have made the dataset publicly available to benefit the broader research community. [^1] We would make our dataset public in the published version
Our investigation first provides clear evidence of the profound structural shift EIP-4844 induced in the market. We observed that as Layer-2s moved their data posting from expensive calldata to cheap blobs, the average block size grew $2.5\times$ to nearly $400$ KB. Simultaneously, the space consumed by traditional transactions from the public mempool shrank dramatically, from over $150$ KB to just $30$ KB. This confirms that builders are actively reallocating block space to accommodate blobs. The fundamental change in builder strategy confirmed their adaptation to the new market. This, in turn, motivated our investigation into its economic efficiency.
However, our investigation reveals that this strategic shift led to widespread economic inefficiency. We identify a significant “crowding out” effect, where builders frequently accepted blob transactions even when their revenue failed to cover the opportunity cost of excluding more profitable transactions readily available in the mempool. This inefficiency was particularly stark in March 2025 (the month following the upgrade), where potential profits from the mempool were over five times greater than the revenue builders actually gained from blobs. Across our entire observation period, we quantify this issue, finding that a striking $29.48\%$ of all blocks containing blobs were unprofitable for this very reason.
The situation for rollups proved to be equally concerning. Our analysis uncovers that their post-EIP-4844 strategies are far from optimal, leading to both direct economic losses from flawed strategies and significant network externalities from blob misuse. Specifically, we identify two prevalent inefficient submission practices. The first, Continuous Sending (accounting for $7\%$ of transactions), occurs when rollups pay redundant fees by sending multiple transactions that could have been bundled into one. The second, Double Sending ($10\%$ of transactions), involves wasting fees on submitting identical, useless blobs after the first one has already been validated. Combined, these self-inflicted inefficiencies have cost rollups a quantifiable loss of at least 186.92 ETH. Furthermore, rollups suffer from the misuse of blob space by external actors. We identify the rise of “Blobscriptions”—a practice of using blobs for on-chain inscriptions rather than their intended purpose of rollup data validation. This creates a negative externality by consuming scarce resources (a maximum of six blobs per block) and congesting the network for legitimate rollups. The impact is significant: during periods of peak activity, this congestion imposed an additional inclusion delay of nearly 19 seconds on legitimate rollup transactions—increasing their average wait time from 24 to 43 seconds—and severely degrading overall service performance.
To address these widespread inefficiencies, we then propose formal solutions. We construct a game-theoretic model to capture the complex interplay between builders and rollups, and from this model, we derive actionable, optimal strategies to correct the identified market failures. For builders, we propose a new pricing framework to ensure provably profitable block construction. Our results also show that, as the number of type-3 transactions increases, the price obtained by builders rises and gradually converges to the revenue that rollups derive from these transactions. For rollups, we find that blockspace constraints do not significantly affect the selection probability of large type-3 transactions. Under these conditions, we then formally prove that bundling multiple blobs into a single transaction is a dominant, cost-saving strategy that maximizes their utility. In essence, our work not only diagnoses the nascent problems of the EIP-4844 market but also provides a clear, evidence-based roadmap toward a more efficient ecosystem.
To summarize, this paper presents the first systematic analysis of the economic dynamics and market strategies on Ethereum following the EIP-4844 upgrade. Our main contributions are:
@@ 1.2. Contribution
    We construct a large-scale dataset to analyze the builder and rollup market in the post-EIP-4844 era. Based on that, our analysis reveals that EIP-4844 successfully altered Ethereum’s transaction structure and significantly reduced data fees.
@@
    We conduct a detailed empirical study of builder and rollup strategies, revealing that both parties operate sub-optimally. Our findings show that builders use inefficient pricing, while rollups exhibit flawed behaviors like redundant blob submissions, resulting in missed revenue and unnecessary costs.
@@
    We develop an optimal pricing model to enable builders to make provably profitable decisions. For rollups, we formally prove that bundling multiple blobs into a single transaction is a dominant, cost-saving strategy.
1.3. Organization
This paper is structured as follows. Section 2 introduces the preliminaries of Layer-2 Rollups, Ethereum Builders, and the details of the EIP-4844 upgrade. Section 3 describes our dataset creation process. Section 4 examines builder strategies in the post-EIP-4844 era by analyzing block composition, revenue, and efficiency, identifying prevalent inefficiencies in their current behavior. Section 5 then evaluates rollup behavior regarding blob and type-3 transaction usage, highlighting their inefficient pricing strategies and the resulting implications. Section 6 develops a game-theoretic model between builders and rollups to derive optimal pricing strategies for builders and utility-maximizing strategies for rollups. Finally, Section 7 reviews related work, and Section 8 concludes the paper.
2. Preliminaries
@@
In this section, we establish the foundational knowledge required to analyze the post-EIP-4844 market. We begin by defining the operational mechanics of key actors, namely builders and rollups. We then delve into the specifics of the EIP-4844 proposal, explaining how data blobs and Type-3 transactions function. A clear grasp of these components is crucial for understanding the subsequent analysis presented in this paper.
2.1. Layer-2 Rollup
Rollups are Layer-2 scaling solutions designed to improve the efficiency and throughput of the Ethereum network by offloading transaction processing from the Ethereum mainnet (Ethereum.org, 2024b). To protect the data availability of Ethereum (Pomerantz, 2024), rollups work by bundling multiple transactions into a single batch. After bundling, the rollup smart contract on Ethereum updates the state with a single, compressed transaction that includes the aggregated data of all individual transactions. This compressed transaction will be submitted to the mainnet, thereby reducing the computational load and gas fee for each individual transaction. Rollups maintain the security and decentralization of Ethereum while providing faster and cheaper transactions, making them a crucial component of the Ethereum scalability roadmap (Smith, 2024).
@@ 2.1. Layer-2 Rollup
2.2. Ethereum Builder
With an upgrade known as the Merge (Foundation., 2023) in 2022, Ethereum has transitioned from a PoW consensus, which is energy intensive, to a PoS mechanism. The Proposer Builder Separation (PBS) mechanism has been proposed as an important development after the Merge (Ethereum.org, 2023). The traditional role of validators, formerly known as miners, is divided into separate entities: MEV-Boost builders and proposers (Buterin, 2021a, b). MEV-Boost builders are responsible for block construction. They select transactions from public mempool (Wackerow, 2024a) or private channels (Wang et al., 2025). The constructed blocks are then forwarded to proposers, accompanied by fee commitments offered by the builders. Proposers endorse the block with the highest fee for the finalization of the blockchain.
Currently, it is optional for Ethereum validators to participate in PBS (Wahrstätter, 2024). Validators who opt not to use PBS and instead build blocks independently are known as Vanilla builders. Vanilla builders receive no private transactions because they are not specialized in block building tasks. They complete both block building and validation tasks simultaneously and currently account for 10% of all builders in the network (Wahrstätter, 2023).
Figure 1 illustrates the current architecture of Ethereum. MEV-boost builders receive transactions from both the public mempool and private channels, whereas Vanilla builders rely solely on the public mempool. Despite differing transaction intake mechanisms, both types of builders contribute to the same outcome — a competitive blockspace market in which users bid for faster inclusion by offering higher priority fees (Liu et al., 2022). Users seeking expedited confirmation increase their priority fees, while those with lower urgency accept longer delays by paying only the base fee.
2.3. EIP-4844 & Type-3 Transaction
@@
Rollups use calldata to store state roots and proofs (Pomerantz, 2024), ensuring that the rollup state could be re-established even if a transaction was challenged. Calldata is the payload contained in a transaction that can be used to store arbitrary information on the chain. While calldata is cost-effective for smaller amounts of data, it is permanently stored in Ethereum state and is therefore relatively expensive for larger data requirements. The high cost of calldata storage becomes a bottleneck for scaling.
@@
In order to improve the scalability and data availability of the network, Ethereum proposed a new sharding design, Danksharding (Ethereum.org, 2024b). Danksharding is a unified sharding design that differs from traditional sharding by managing data storage and availability in a more streamlined way. Instead of increasing the space for transactions, Ethereum sharding provides more space for blobs of data, which the Ethereum protocol itself does not interpret.
EIP-4844, also known as Proto-Danksharding, implements the same transaction formats and validation rules. EIP-4844 was launched in March as a core part of the Dencun upgrade in 2024 (Ethereum.org, 2024a), thereby facilitating the network’s preparation for full Danksharding by allowing the inclusion of large data in transactions (Pomerantz, 2024). The EIP-4844 represents a significant milestone in Ethereum’s roadmap (Ethereum.org, 2024c), aiming to improve the scalability, security, and usability of the network.
@@
There were three types of transaction before EIP-4844: type-0, type-1 and type-2 transaction. Note, type-2 transactions always bring a priority fee for inclusion by builders quickly. EIP-4844 implements on a new type: type-3 transactions (Park et al., 2024). These are similar to regular type-2 transactions, but carry an additional data component called a ‘blob’. A blob is a large, temporary data payload attached to a type-3 transaction. Unlike calldata, blob data is not stored permanently on-chain. Instead, only a commitment to the blob is included on-chain, significantly reducing storage costs (Pomerantz, 2024). These oversized blobs (128 KB) are much cheaper than the equivalent amount of calldata because they are not permanently stored on Ethereum (Pomerantz, 2024), and rollups can also send their type-3 transactions from public mempool or private channels. This new type of transaction meets the growing need for efficient data processing. The blob data remains inaccessible during the execution of the Ethereum virtual machine, and the validator can only see the commitment to the blob, but these blobs can be used for Layer-2 scaling solutions and other off-chain protocols that require large amounts of data throughput (Gogol et al., 2024a).
3. Data Collection Methodology